OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 22 and the last term is 1,299,078?

OpenStudy (anonymous):

1,395,030 1,461,460 1,527,890 1,594,320

OpenStudy (anonymous):

@mathmate

OpenStudy (abmon98):

\[S(n)=a(1-r^n)/(1-r)\] \[ar ^{n-1}\] \[ar^5\] \[22r^5=1,299,078\]\[r=(1,299,078/22)^{1/5}=9\] \[S(6)=22(1-9^6)/(1-9)\]

OpenStudy (mathmate):

@Abmon98 has shown all the work. There is nothing left to do except use the calculator to find the numerical answer, which is one of the choices.